# The physical and chemical processes in protoplanetary disks: constraints on the composition of comets

### v3

@@ 3.1.1 CO and HO
In molecular clouds, CO and HO ice are major Carbon and Oxygen carriers with an abundance of relative to H, and thus are naively expected so in the disks. While HO snowlines are too close ( several au) to the central star to be spatially resolved in the observations, CO snowline is at au around T Tauri stars (§2.2). The observation of CO snowline is, however, not straightforward, since CO gas is abundant in the warm molecular layer even outside the CO snowline. NH line is used to trace the CO snowline; its distribution anti-correlates with CO, since it is destroyed (NH + CO N + HCO) when CO is abundant in the gas phase (Qi et al., 2013a, 2019; Aikawa et al., 2015). Alternative tracers are rare isotopes of CO (e.g. CO), whose emission is expected to be optically thin (Zhang et al., 2017). In TW Hya, for example, the CO snowline is estimated to be au and 20.5 au from NH and CO, respectively (Qi et al., 2013a; Zhang et al., 2017).
@@
The observations of rare isotopes of CO also suggest that the gaseous abundance of CO is actually lower than even in the warm molecular layer. The CO column densities derived from the observations are compared with the disk models which are constrained by dust continuum emission. As an example, Fig. 5 shows the radial distribution of CO and dust column densities derived from CO () and dust continuum (Zhang et al., 2021; Sierra et al., 2021). Assuming the gas/dust mass ratio is 100, the CO depletion factor, i.e. the ratio of canonical CO abundance () divided by the estimated CO abundance in the warm molecular layer, varies from 1 to 100 among disks in radii outside the CO snowline (Miotello et al., 2017; Zhang et al., 2019, 2021). While the CO abundance tends to increase inwards, the CO depletion factor is even inside the CO snowline in some disks. An alternative interpretation of these observations is that the gas/dust mass ratio is lower than 100. The submillimeter HD emission lines, which are observed by Herschel towards a limited number of disks, suggest CO depletion rather than gas depletion (Bergin et al., 2013; Favre et al., 2013; McClure et al., 2016). It should however be noted that the evaluation of disk gas mass from HD emission also depends on the assumed thermal structure, i.e. disk models (Miotello et al., 2023).
@@
Herschel also observed a few HO lines in several disks (Hogerheijde et al., 2011; Salinas et al., 2016; Du et al., 2017). While the beam size (e.g. 37 at 557 GHz) is too large to provide any information on the spatial distribution, the emission lines of HO are significantly fainter than the predictions of disk models, in which HO is photodesorbed from ice outside its snowline. It suggests that HO ice-coated grains are mostly sedimented to the disk midplane where photodesorption is inefficient, and could even be radially drifted inwards (§2.3). The low CO abundance in the warm molecular layer can be explained similarly by the cold-finger effect; while CO is in the gas phase in the warm layer, it could be brought to the cold midplane by weak turbulence to freeze out onto grains. Conversion of CO to less volatile molecules (e.g. CO ice and CHOH ice) would also help (e.g. Furuya and Aikawa, 2014; Kama et al., 2016; Krijt et al., 2020; Furuya et al., 2022a).
@@
It is noteworthy that HO ice and CO are expected to be the dominant reservoirs of oxygen and carbon in the initial condition (i.e. gas and solids accreted from the molecular clouds). While both can be depleted from the warm molecular layer as described above, lower volatility of HO would result in heavier depletion of oxygen than carbon, and an elemental abundance of C/O in the warm molecular layer. The high C/O ratio is supported by the observations of other molecules as explained below.
@@
3.1.2 Hydrocarbons: CH and -CH
Unsaturated hydrocarbons, CH and -CH, are expected to be abundant in the surface PDR layer, which is supported by the imaging observation of edge-on disks (Ruíz-Rodríguez et al., 2021) and their relatively high excitation temperature (20 K 50 K) (Guzmán et al., 2021; Cleeves et al., 2021).
The abundance of CH is also sensitive to the C/O ratio in the gas phase; the comparison of its line brightness between observations and models indicates C/O (Bergin et al., 2016; Bergner et al., 2019; Miotello et al., 2019). However, variation among disks should be noted; some disks are faint in CH emission (Miotello et al., 2019). Spatially resolved observations show that the CH emissions often show a ring-like structure. Some rings coincide with the prominent gaps seen in the dust continuum, which is consistent with the expectation that CH abundance is enhanced by UV penetration. But not all dust gap correlates with CH ring, and vice visa (Guzmán et al., 2021; Law et al., 2021a).
-CH is detected in several disks, although its emission lines are weaker than those of CH (Qi et al., 2013b). The column density ratio of -CH/CH is relatively constant with in four disks where this ratio is derived (Ilee et al., 2021) (Fig. 5). These observations and theoretical models suggest that -CH is formed in the gas phase together with CH.
@@ 3.1.3 HCN and CN
CN is also predicted to be abundant in the surface PDR layer, which is supported by the excitation analysis (Teague and Loomis, 2020; Bergner et al., 2021). A ring-like distribution of CN emission is found in some disks, which is explained by CN formation via the reaction of N atom with UV-pumped vibrationally excited H (Cazzoletti et al., 2018b). HCN, on the other hand, is more susceptible to photodissociation than CN; it is dissociated by UV at longer wavelengths than for CN, and by Ly emitted from T Tauri stars.
@@ 3.1.4 S-bearing molecules
As in molecular clouds, the abundances of gaseous S-bearing molecules in the disks are much lower than expected from model predictions assuming the Solar elemental abundance, which suggests significant depletion of sulfur in the solid phase (see §4.2 and the chapter by Bergin et al.). Six S-bearing molecules are detected in disks: CS, SO, HS, HCS, CS and SO. Among them, CS is the most readily detected, and its column density is estimated to be cm in several disks (Le Gal et al., 2019). Detection of other S-bearing molecules is rather limited. SO is first detected in AB Aur disk by Fuente et al. (2010). Guilloteau et al. (2016) observed SO in 20 disks using IRAM 30 m and detected the emission only in 4 disks. Prevalence of CS over SO may suggest a high C/O ratio in disks (Le Gal et al., 2021). HS, the dominant sulfur-bearing molecule in comets, is first detected in disks by Phuong et al. (2018), and is recently detected in 4 additional disks (Rivière-Marichalar et al., 2021). HCS is detected in the disk of Herbig Ae star, MWC 480 (Le Gal et al., 2019).
Recently SO was first detected in the disk of IRS 48, which is a transitional disk (i.e. a disk with a central hole in dust continuum emission) with an asymmetric dust emission peak (Booth et al., 2021a). The emission peak would correspond to the local pressure maximum, at which dust grains are trapped (§2). Both SO and SO are bright only at this dust peak, while CS is not detected in this disk. The low CS/SO ratio indicates a low C/O ratio () at the dust trap of IRS 48.
@@ 3.1.5 Complex Organic Molecules
In the astrochemistry community, organic molecules consist of 6 atoms and more are called complex organic molecules (COMs). Methanol, the most abundant and prototypical COM in molecular clouds, is so far detected only in TW Hya, HD 100546, and IRS 48 (Walsh et al., 2016; Booth et al., 2021b; van der Marel et al., 2021) in spite of deep searches in some other disks (Loomis et al., 2020; Carney et al., 2019). It is also the key molecule to investigate the relation between ice and gas in disks; since its formation in the gas phase is known to be inefficient, gaseous CHOH should be mainly desorbed from ice, which in turn, is formed by hydrogenation of CO on cold ( K) ice surface. The CHOH emission in TW Hya is very weak and is in a ring region outside the CO snowline ( au) and inside the millimeter dust continuum edge. Since the dust temperature in the emitting region is below its sublimation temperature ( K), CHOH is considered to be desorbed by non-thermal processes. HD 100546, on the other hand, is a warm Herbig Ae disk with a central hole in the dust continuum. Most of the CHOH emission originates from the spatially unresolved central region ( au) and thus could be tracing the thermally sublimated CHOH. The warm dust temperature ( K) in the disk of HD 100546 indicates that CHOH is not formed in the disk, but inherited from molecular clouds (Booth et al., 2021b). In IRS 48, the CHOH emission spatially coincides with the dust trap and shows a high excitation temperature K. Gaseous methanol abundance could be enhanced by a combination of ice-coated pebble concentration in the dust trap and irradiation heating at the edge of the central hole (van der Marel et al., 2021). Recently, CHOCH and CHOCHO are also detected at the same position as CHOH in IRS 48 (Brunken et al., 2022).
@@
Formic acid (HCOOH), which is often detected in the protostellar cores with bright CHOH emissions, is so far detected only in TW Hya (Favre et al., 2018). HCO is another relevant molecule to CHOH. It is an intermediate product of CO hydrogenation to form CHOH on grain surfaces, while it can be formed via gas-phase reactions as well. HCO is also considered to be a precursor species of organics in meteorites (Cody et al., 2011). Pegues et al. (2020) detected HCO in 13 disks out of 15, and derived an excitation temperature of K and column density of cm. The second bottom panels in Fig. 5 show the radial distribution of HCO in 3 T Tauri disks assuming the excitation temperature of 20 K (Guzmán et al., 2021).
@@
A nitrogen bearing COM, CHCN, and another nitrile species HCN are detected in several disks (Öberg et al., 2015; Bergner et al., 2018). Ilee et al. (2021) spatially resolved the emission lines of these nitriles in disks of GM Aur, AS 209, HD 163296, and MWC 480; the emission is compact ( au) comparable to the extent of the dust continuum. Unlike CHOH, these species can be formed in the gas phase. The weak correlation of their emission distributions with the dust continuum, however, may indicate the role of ice reservoir for their formation. The excitation temperatures ( K) are lower than their sublimation temperatures, indicating that they are non-thermally desorbed from ice or formed in the gas phase from the species with lower sublimation temperatures. Their column densities ( cm) (Fig. 5) are higher than predicted in static disk models; either or a combination of grain growth and turbulent mixing would enhance their abundances (Semenov and Wiebe, 2011; Furuya and Aikawa, 2014; Öberg et al., 2015).
@@ 3.2 Molecules in warm and hot gas
Abundant warm molecular gas (K) is generally observed in typical protoplanetary disks surrounding stars of widely different masses (Brittain et al., 2003; Blake and Boogert, 2004; Brown et al., 2013). The regions of protoplanetary disks where such warm gas is present are often referred to as the “inner disk”. The inner disk may similarly be defined as corresponding to disk radii inside the water snowline. While refractory carbon and silicate-dominated dust sublimate at 1500 K, i.e. at the radius of 0.01-0.1 au for typical protoplanetary disks, many chemically important volatile molecular species (water, CO, CHOH, CH, NH, etc.) persist in the gas phase down to dust temperatures of 100 K or below. Such temperatures are achieved at 1-2 au in the midplane for disks around young solar-mass stars, and at distances of up to 10 au in the superheated disk surface (Sasselov and Lecar, 2000; Garaud and Lin, 2007) (Fig, 3). Indeed, high-resolution spectroscopy and spectral imaging have demonstrated that the emission originates well inside their respective snowlines (Goto et al., 2006; Pontoppidan et al., 2008), and that the location of the snowline likely varies with the evolutionary stage of the star-disk system (Hsieh et al., 2019, see also §2.2).
@@
The relative abundances of warm gas-phase volatiles in inner disk surfaces appear to be different from those observed in ices in the cold interstellar medium, and in Solar system comets. This indicates that the observed inner disk chemistry is not directly inherited from cold ice chemistry, but is significantly altered by local processes. Concurrent measurements of H and CO using UV spectroscopy suggest that inner disk CO abundances range from slightly depleted up to canonical ([CO/H]10–10) (France et al., 2014; Cauley et al., 2021). Other observable species, such as NH, CH, and CO, are further depleted relative to CO by orders of magnitude, clearly inconsistent with primordial ice chemistry (Mandell et al., 2012; Pontoppidan and Blevins, 2014; Bosman et al., 2017; Pontoppidan et al., 2019b), but consistent with current predictions for warm gas-phase chemistry, which tends to destroy particularly those species, driving nitrogen into N and carbon into CO (Agúndez et al., 2008; Walsh et al., 2015). Water, on the other hand, is commonly highly abundant in disks around low-mass and solar-mass young stars, as revealed by an extensive sample of disks observed by Spitzer (Carr and Najita, 2008; Salyk et al., 2008; Pontoppidan et al., 2010; Carr and Najita, 2011), with retrieved abundances consistent with [CO/HO]1 (Salyk et al., 2011). JWST is currently confirming the ubiquitous presence of mid-infrared emission from warm gas in inner disks around young stars of all masses, and demonstrate a widening diversity of relative molecular abundances from water and organics (Tabone et al., 2023; Banzatti et al., 2023; van Dishoeck et al., 2023).

### v4

@@ 3.1.1 CO and H2O
In molecular clouds, CO and H2O ice are major Carbon and Oxygen carriers with an abundance of relative to H2, and thus are naively expected so in the disks. While H2O snowlines are too close ( several au) to the central star to be spatially resolved in the observations, CO snowline is at au around T Tauri stars (§2.2). The observation of CO snowline is, however, not straightforward, since CO gas is abundant in the warm molecular layer even outside the CO snowline. N2H+ line is used to trace the CO snowline; its distribution anti-correlates with CO, since it is destroyed (N2H+ + CO N2 + HCO+) when CO is abundant in the gas phase (Qi et al., 2013a, 2019; Aikawa et al., 2015). Alternative tracers are rare isotopes of CO (e.g. 13C18O), whose emission is expected to be optically thin (Zhang et al., 2017). In TW Hya, for example, the CO snowline is estimated to be au and 20.5 au from N2H+ and 13C18O, respectively (Qi et al., 2013a; Zhang et al., 2017).
@@
The observations of rare isotopes of CO also suggest that the gaseous abundance of CO is actually lower than even in the warm molecular layer. The CO column densities derived from the observations are compared with the disk models which are constrained by dust continuum emission. As an example, Fig. 5 shows the radial distribution of CO and dust column densities derived from C18O () and dust continuum (Zhang et al., 2021; Sierra et al., 2021). Assuming the gas/dust mass ratio is 100, the CO depletion factor, i.e. the ratio of canonical CO abundance () divided by the estimated CO abundance in the warm molecular layer, varies from 1 to 100 among disks in radii outside the CO snowline (Miotello et al., 2017; Zhang et al., 2019, 2021). While the CO abundance tends to increase inwards, the CO depletion factor is even inside the CO snowline in some disks. An alternative interpretation of these observations is that the gas/dust mass ratio is lower than 100. The submillimeter HD emission lines, which are observed by Herschel towards a limited number of disks, suggest CO depletion rather than gas depletion (Bergin et al., 2013; Favre et al., 2013; McClure et al., 2016). It should however be noted that the evaluation of disk gas mass from HD emission also depends on the assumed thermal structure, i.e. disk models (Miotello et al., 2023).
@@
Herschel also observed a few H2O lines in several disks (Hogerheijde et al., 2011; Salinas et al., 2016; Du et al., 2017). While the beam size (e.g. 37 at 557 GHz) is too large to provide any information on the spatial distribution, the emission lines of H2O are significantly fainter than the predictions of disk models, in which H2O is photodesorbed from ice outside its snowline. It suggests that H2O ice-coated grains are mostly sedimented to the disk midplane where photodesorption is inefficient, and could even be radially drifted inwards (§2.3). The low CO abundance in the warm molecular layer can be explained similarly by the cold-finger effect; while CO is in the gas phase in the warm layer, it could be brought to the cold midplane by weak turbulence to freeze out onto grains. Conversion of CO to less volatile molecules (e.g. CO2 ice and CH3OH ice) would also help (e.g. Furuya and Aikawa, 2014; Kama et al., 2016; Krijt et al., 2020; Furuya et al., 2022a).
@@
It is noteworthy that H2O ice and CO are expected to be the dominant reservoirs of oxygen and carbon in the initial condition (i.e. gas and solids accreted from the molecular clouds). While both can be depleted from the warm molecular layer as described above, lower volatility of H2O would result in heavier depletion of oxygen than carbon, and an elemental abundance of C/O in the warm molecular layer. The high C/O ratio is supported by the observations of other molecules as explained below.
@@
3.1.2 Hydrocarbons: C2H and -C3H2
Unsaturated hydrocarbons, C2H and -C3H2, are expected to be abundant in the surface PDR layer, which is supported by the imaging observation of edge-on disks (Ruíz-Rodríguez et al., 2021) and their relatively high excitation temperature (20 K 50 K) (Guzmán et al., 2021; Cleeves et al., 2021).
The abundance of C2H is also sensitive to the C/O ratio in the gas phase; the comparison of its line brightness between observations and models indicates C/O (Bergin et al., 2016; Bergner et al., 2019; Miotello et al., 2019). However, variation among disks should be noted; some disks are faint in C2H emission (Miotello et al., 2019). Spatially resolved observations show that the C2H emissions often show a ring-like structure. Some rings coincide with the prominent gaps seen in the dust continuum, which is consistent with the expectation that C2H abundance is enhanced by UV penetration. But not all dust gap correlates with C2H ring, and vice visa (Guzmán et al., 2021; Law et al., 2021a).
-C3H2 is detected in several disks, although its emission lines are weaker than those of C2H (Qi et al., 2013b). The column density ratio of -C3H2/C2H is relatively constant with in four disks where this ratio is derived (Ilee et al., 2021) (Fig. 5). These observations and theoretical models suggest that -C3H2 is formed in the gas phase together with C2H.
@@ 3.1.3 HCN and CN
CN is also predicted to be abundant in the surface PDR layer, which is supported by the excitation analysis (Teague and Loomis, 2020; Bergner et al., 2021). A ring-like distribution of CN emission is found in some disks, which is explained by CN formation via the reaction of N atom with UV-pumped vibrationally excited H2 (Cazzoletti et al., 2018b). HCN, on the other hand, is more susceptible to photodissociation than CN; it is dissociated by UV at longer wavelengths than for CN, and by Ly emitted from T Tauri stars.
@@ 3.1.4 S-bearing molecules
As in molecular clouds, the abundances of gaseous S-bearing molecules in the disks are much lower than expected from model predictions assuming the Solar elemental abundance, which suggests significant depletion of sulfur from the gas to the solid phase (see §4.2 and the chapter by Bergin et al.). Six S-bearing molecules are detected in disks: CS, SO, H2S, H2CS, C2S and SO2. Among them, CS is the most readily detected, and its column density is estimated to be cm-2 in several disks (Le Gal et al., 2019). Detection of other S-bearing molecules is rather limited. SO is first detected in AB Aur disk by Fuente et al. (2010). Guilloteau et al. (2016) observed SO in 20 disks using IRAM 30 m and detected the emission only in 4 disks. Prevalence of CS over SO may suggest a high C/O ratio in disks (Le Gal et al., 2021). H2S, the dominant sulfur-bearing molecule in comets, is first detected in disks by Phuong et al. (2018), and is recently detected in 4 additional disks (Rivière-Marichalar et al., 2021). H2CS is detected in the disk of Herbig Ae star, MWC 480 (Le Gal et al., 2019).
Recently SO2 was first detected in the disk of IRS 48, which is a transitional disk (i.e. a disk with a central hole in dust continuum emission) with an asymmetric dust emission peak (Booth et al., 2021a). The emission peak would correspond to the local pressure maximum, at which dust grains are trapped (§2). Both SO and SO2 are bright only at this dust peak, while CS is not detected in this disk. The low CS/SO ratio indicates a low C/O ratio () at the dust trap of IRS 48.
@@ 3.1.5 Complex Organic Molecules
In the astrochemistry community, organic molecules consist of 6 atoms and more are called complex organic molecules (COMs). Methanol, the most abundant and prototypical COM in molecular clouds, is so far detected only in TW Hya, HD 100546, IRS 48, and HD 169142 (Walsh et al., 2016; Booth et al., 2021b; van der Marel et al., 2021; Booth et al., 2023) in spite of deep searches in some other disks (Loomis et al., 2020; Carney et al., 2019). It is also the key molecule to investigate the relation between ice and gas in disks; since its formation in the gas phase is known to be inefficient, gaseous CH3OH should be mainly desorbed from ice, which in turn, is formed by hydrogenation of CO on cold ( K) ice surface. The CH3OH emission in TW Hya is very weak and is in a ring region outside the CO snowline ( au) and inside the millimeter dust continuum edge. Since the dust temperature in the emitting region is below its sublimation temperature ( K), CH3OH is considered to be desorbed by non-thermal processes. HD 100546, on the other hand, is a warm Herbig Ae disk with a central hole in the dust continuum. Most of the CH3OH emission originates from the spatially unresolved central region ( au) and thus could be tracing the thermally sublimated CH3OH. The warm dust temperature ( K) in the disk of HD 100546 indicates that CH3OH is not formed in the disk, but inherited from molecular clouds (Booth et al., 2021b). In IRS 48, the CH3OH emission spatially coincides with the dust trap and shows a high excitation temperature K. Gaseous methanol abundance could be enhanced by a combination of ice-coated pebble concentration in the dust trap and irradiation heating at the edge of the central hole (van der Marel et al., 2021). Recently, CH3OCH3 and CH3OCHO are also detected at the same position as CH3OH in IRS 48 (Brunken et al., 2022).
@@
Formic acid (HCOOH), which is often detected in the protostellar cores with bright CH3OH emissions, is so far detected only in TW Hya (Favre et al., 2018). H2CO is another relevant molecule to CH3OH. It is an intermediate product of CO hydrogenation to form CH3OH on grain surfaces, while it can be formed via gas-phase reactions as well. H2CO is also considered to be a precursor species of organics in meteorites (Cody et al., 2011). Pegues et al. (2020) detected H2CO in 13 disks out of 15, and derived an excitation temperature of K and column density of cm-2. The second bottom panels in Fig. 5 show the radial distribution of H2CO in 3 T Tauri disks assuming the excitation temperature of 20 K (Guzmán et al., 2021).
@@
A nitrogen bearing COM, CH3CN, and another nitrile species HC3N are detected in several disks (Öberg et al., 2015; Bergner et al., 2018). Ilee et al. (2021) spatially resolved the emission lines of these nitriles in disks of GM Aur, AS 209, HD 163296, and MWC 480; the emission is compact ( au) comparable to the extent of the dust continuum. Unlike CH3OH, these species can be formed in the gas phase. The weak correlation of their emission distributions with the dust continuum, however, may indicate the role of ice reservoir for their formation. The excitation temperatures ( K) are lower than their sublimation temperatures, indicating that they are non-thermally desorbed from ice or formed in the gas phase from the species with lower sublimation temperatures. Their column densities ( cm-2) (Fig. 5) are higher than predicted in static disk models; either or a combination of grain growth and turbulent mixing would enhance their abundances (Semenov and Wiebe, 2011; Furuya and Aikawa, 2014; Öberg et al., 2015).
@@ 3.2 Molecules in warm and hot gas
Abundant warm molecular gas (K) is generally observed in typical protoplanetary disks surrounding stars of widely different masses (Brittain et al., 2003; Blake and Boogert, 2004; Brown et al., 2013). The regions of protoplanetary disks where such warm gas is present are often referred to as the “inner disk”. The inner disk may similarly be defined as corresponding to disk radii inside the water snowline. While refractory carbon and silicate-dominated dust sublimate at 1500 K, i.e. at the radius of 0.01-0.1 au for typical protoplanetary disks, many chemically important volatile molecular species (water, CO2, CH3OH, CH4, NH3, etc.) persist in the gas phase down to dust temperatures of 100 K or below. Such temperatures are achieved at 1-2 au in the midplane for disks around young solar-mass stars, and at distances of up to 10 au in the superheated disk surface (Sasselov and Lecar, 2000; Garaud and Lin, 2007) (Fig, 3). Indeed, high-resolution spectroscopy and spectral imaging have demonstrated that the emission originates well inside their respective snowlines (Goto et al., 2006; Pontoppidan et al., 2008), and that the location of the snowline likely varies with the evolutionary stage of the star-disk system (Hsieh et al., 2019, see also §2.2).
@@
The relative abundances of warm gas-phase volatiles in inner disk surfaces appear to be different from those observed in ices in the cold interstellar medium, and in Solar system comets. This indicates that the observed inner disk chemistry is not directly inherited from cold ice chemistry, but is significantly altered by local processes. Concurrent measurements of H2 and CO using UV spectroscopy suggest that inner disk CO abundances range from slightly depleted up to canonical ([CO/H2]10-5–10-4) (France et al., 2014; Cauley et al., 2021). Other observable species, such as NH3, CH4, and CO2, are further depleted relative to CO by orders of magnitude, clearly inconsistent with primordial ice chemistry (Mandell et al., 2012; Pontoppidan and Blevins, 2014; Bosman et al., 2017; Pontoppidan et al., 2019b), but consistent with current predictions for warm gas-phase chemistry, which tends to destroy particularly those species, driving nitrogen into N2 and carbon into CO (Agúndez et al., 2008; Walsh et al., 2015). Water, on the other hand, is commonly highly abundant in disks around low-mass and solar-mass young stars, as revealed by an extensive sample of disks observed by Spitzer (Carr and Najita, 2008; Salyk et al., 2008; Pontoppidan et al., 2010; Carr and Najita, 2011), with retrieved abundances consistent with [CO/H2O]1 (Salyk et al., 2011). JWST is currently confirming the ubiquitous presence of mid-infrared emission from warm gas in inner disks around young stars of all masses, and demonstrate a widening diversity of relative molecular abundances from water and organics (Tabone et al., 2023; Banzatti et al., 2023; van Dishoeck et al., 2023).